A body slides down a rough inclined plane of angle 30° to the horizontal.find the distance covered after 15s from rest, assuming g=9.8m/s²


1
Expert's answer
2021-07-07T11:37:33-0400

Gives

"\theta=30°"

t=15sec

g=9.8m/sec"^2"

"F=mgsin\theta"

"ma=mgsin\theta"

"a=gsin\theta"

"s=ut+\frac{1}{2}at^2"

"s=0+\frac{1}{2}(gsin\theta)t^2"


"s=\frac{1}{2}\times9.8\times sin30°\times15^2=551.25m"